a narrow glass tube 80 cm long opens at both the ends is half immersed in Hg . now top is closed and taken out of Hg a 20cm long Hg column remain in tube . find atmospheric pressure.
 

Grade:12

2 Answers

Kanishk goyal
7 years ago
When the glass tube was in mercury with half dipped Pressure on top of mecury should be equal to mercury pressure Therefore that pressure will be P(1) =13.6g(40)Now after taking tube out thumb is placed on the top so This should apply there also there now in this condition the pressure of mercury + P1= atmospheric pressure applying now on the other open end ThereforeAtmospheric pressure=P1+13.6g(20) = 13.6g(40)+13.6g(20) =13.6g(60)Atmospheric pressure =60 cm of Hg column
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

When the glass tube was in mercury with half dipped Pressure on top of mecury should be equal to mercury pressure
Therefore that pressure will be P(1) =13.6g(40)
Now after taking tube out thumb is placed on the top so
This should apply there also there now in this condition the pressure of mercury + P1= atmospheric pressure applying now on the other open end
ThereforeAtmospheric pressure = P1+13.6g(20) = 13.6g(40)+13.6g(20) =13.6g(60)
Atmospheric pressure =60 cm of Hg column

Thanks and Regards